OpenStudy (anonymous):

Calculus help with integration and finding volumes? A solid lies between the planes perpendicular to the x-axis at x = sqrt(3) and x = -sqrt(3) The cross-sections perpendicular to the x-axis are circles whose diameters stretch from y=4/sqrt(1+x^2) to y=-4/sqrt(1+x^2) Find the volume of the solid in cubic units. Step by step instructions would be wonderfully helpful!

OpenStudy (anonymous):

OK. Let's walk through this.

OpenStudy (anonymous):

Sure! I believe this would be the disk method, so you would integrate piR^2. I am not sure about what to do with the + and - y equations as R, they are the same except for their signs. I also do not understand how I would integrate (4/sqrt(1+x^2))^2 - I know that if it was not squared this would fit the arctan(x) integration rule, but it is squared so I am not sure what to do.

OpenStudy (anonymous):

So I would just integrate \[4\div(\sqrt{1+x ^{2}}\] then? I'm still not entirely sure how to do this.

OpenStudy (anonymous):

Sorry, I re-read the question, you're right.

OpenStudy (anonymous):

Is this for single-variable or multi-variable? For single you do need to do it with cross-sections, for multivar you can do it with a double or triple integral.

OpenStudy (anonymous):

I think single? I haven't heard of double or triple integrals. Also, if this helps, the answer is 32pi^2/3. I just would like to know how to get there.

OpenStudy (anonymous):

Yeah, this is for single. So we're going to add up a lot of circles to get our shape.

OpenStudy (anonymous):

We know that the radius of each circle is \[\frac{4}{\sqrt{1+x^2}}\]so the area of each circle is \[\frac{16\pi}{1+x^2}\]right?

OpenStudy (anonymous):

Okay yes I see how that would happen.

OpenStudy (anonymous):

So now we're just going to integrate that from -\sqrt{3} to \sqrt{3} (this is basically the disk method, but you're really only supposed to do that with solids of rotation).

OpenStudy (anonymous):

\[\int^\sqrt{3}_{-\sqrt{3}} \frac{16\pi}{1+x^2} dx=16\pi \int^\sqrt{3}_{-\sqrt{3}} \frac{dx}{1+x^2} =16\pi\left[ \arctan(x) \right]^\sqrt{3}_{-\sqrt{3}}=16\pi\left(\frac{\pi}{3}-\frac{-\pi}{3}\right)=\frac{32\pi}{3} \]Sorry for the delay, my LaTeX skills are subpar.

OpenStudy (anonymous):

Okay, that looks pretty good except for the fact that pi is supposed to be squared. Where would that happen?

OpenStudy (anonymous):

It is. To continue that line which got cut off, \[16\pi\left(\frac{\pi}{3}-\frac{-\pi}{3}\right)=16\pi\frac{2\pi}{3}=\frac{32\pi^2}{3}\]

OpenStudy (anonymous):

Thank you!